How do I solve x^2+(x+3)^2=12^2?
Answer by ankor@dixie-net.com(22740)

How do I solve x^2 + (x+3)^2 = 12^2
FOIL(x+3)(x+3), square 12
x^2 + x^2 + 6x + 9 = 144
combine like terms on the left to form a quadratic equation
2x^2 + 6x + 9 - 144 = 0
2x^2 + 6x - 135 = 0
we have to use the quadratic equation to find x here
I got a positive solution: x = 6.85..